# Practical Challenges in Executing Shor’s Algorithm on Existing Quantum Platforms

### v2

@@ 2 Understanding Different Quantum Computers
2.1 Quantum Computers with Synthetic Qubits
@@
2.2 Quantum Computers with Natural Qubits
@@ 3 Review on Shor’s Implementations
3.1 Shor’s Algorithm
As discussed in [5], quantum computers excel at solving problems with modest classical input/output data but high computational complexity. Shor’s algorithm [1, 2] is a canonical example: it factors an $N$-bit integer in time polynomial in $N$, in stark contrast to the sub-exponential but super-polynomial complexity of the best-known classical factoring algorithms.
@@ 3.1 Shor’s Algorithm
    Find the period (quantum). Use a QFT–based QPE routine to infer $r$ from the periodic structure of $a^{x}\bmod N$. This order-finding step is classically believed to require exponential time in the bit-length of $N$, but can be solved in polynomial time on a fault-tolerant quantum computer.
@@
3.2 Shor’s Implementation: State-of-the-Art
@@ 4.1 Setup and Notation
Let $N$ be the composite to factor, $a$ a random coprime modulo $N$, and $r$ the order of $a\bmod N$. The phase register in QPE has $t$ qubits, so its grid has size $L=2^{t}$ and the measured outcome is $y\in\{0,\dots,L-1\}$. In the ideal case, QPE produces peaks near the rationals $s/r$ for $s=0,\dots,r-1$. We choose:
$t\;\geq\;2\left\lceil\log_{2}N\right\rceil$ (2)
@@
However, with modern post-processing (Ekerå et al. [24]), the required phase precision can be cut to roughly $t\approx n+O(\log n)$ while maintaining high single-run success via limited classical searches, thereby reducing depth and repetitions [24, 25]. Thus, we allow ourselves to use a lower $t=10<12$ for the case $N=35$ while still considering the framework scalable  [^2] Further details on the experimental certification method and refinements regarding scalability in the context of Shor’s algorithm are provided in the Appendix A.1..
4.2 Public Quantum Platforms
@@ 4.3 Results (N=15, 21, 35)
We then executed the same circuits on ibm_torino, after transpilation to the device topology. For each modulus $N$, we focus on detecting a statistically significant excess of probability mass in the theoretically predicted windows around $sL/r$ relative to the uniform baseline. The formulas of the uniform baseline and the empirical probability used for the experiments, rigorously derived in the Appendix A.1, are the following:
$b=\frac{|acceptance\_bins|\times(2\omega_{0}+1)}{L}\text{ and }\hat{p}=\frac{hits}{shots}$
We consider three composite numbers: $N=15$, $N=21$, and $N=35$. In the Appendix subsection A.7, we further discuss the shape of the empirical histograms compared to the ideal noiseless expected distributions.
@@ 5 Limitations and Challenges
Our experiments focus primarily on detecting a statistically significant quantum signal in QPE histograms, rather than fully establishing polylogarithmic scaling of the expected runtime for a fixed acceptance parameter $k$ (see Appendix A.6). For each small modulus $N$, we tune window sizes and shot counts to test whether the mass in acceptance regions exceeds a uniform baseline. This is sufficient to detect coherence in the presence of noise but does not yet constitute a full scalability study.
@@ Acknowledgments
We gratefully acknowledge Evgueni Rousselot for his initial experiments, which were instrumental in cenabling and motivating the investigations further developed in this paper.
@@ A.2 Per-Run Success Probability
A.3 Importance of the Quantum Signal
@@
A.4 Certifying Substantial Quantum Signal in Practice
@@
A.5 Shor in the Ideal Regime
@@
A.6 Shor in the Experimental Regime
@@
A.7 Effect of Decoherence on the QPE Output Distribution

### v3

@@ 2 Understanding Different Quantum Computers
2.A Quantum Computers with Synthetic Qubits
@@
2.B Quantum Computers with Natural Qubits
@@ 3 Review on Shor’s Implementations
3.A Shor’s Algorithm
As discussed in [5], quantum computers excel at solving problems with modest classical input/output data but high computational complexity. Shor’s algorithm [1, 2] is a canonical example: it factors an $n$-bit integer $N$ in time polynomial in $n$, in stark contrast to the sub-exponential but super-polynomial complexity of the best-known classical factoring algorithms.
@@ 3.A Shor’s Algorithm
    Find the period (quantum). Use a QFT–based QPE routine to infer $r$ from the periodic structure of $a^{x}\bmod N$.Classically, order finding in $\mathbb{Z}_{N}^{\times}$ is not known to admit a polynomial-time algorithm and is believed to require sub-exponential time in the bit-length of $N$. In contrast, a fault-tolerant quantum computer can solve this problem in polynomial time.
@@
3.B Shor’s Implementation: State-of-the-Art
@@ 4.A Setup and Notation
Let $N$ be the composite to factor, $a$ a random coprime modulo $N$, and $r$ the order of $a\bmod N$. The phase register in QPE has $t$ qubits, so its grid has size $L=2^{t}$ and the measured outcome is $y\in\{0,\dots,L-1\}$. In the ideal case, QPE produces peaks near the rationals $s/r$ for $s=0,\dots,r-1$. We choose, denoting $n=\lceil\log_{2}N\rceil$ the number of bits,
$t\;\geq\;2n$ (2)
@@
However, with modern post-processing (Ekerå  [24, 25]), the required phase precision can be cut to roughly $t\approx n+O(\log n)$ while maintaining high single-run success via limited classical searches, thereby reducing depth and repetitions. Thus, we allow ourselves to use a lower $t=10<12$ for the case $N=35$ while still considering the framework scalable  [^2] Further details on the experimental certification method and refinements regarding scalability in the context of Shor’s algorithm are provided in the Appendix A.A..
4.B Public Quantum Platforms
@@ 4.C Results (N=15, 21, 35)
We then executed the same circuits on ibm_torino, after transpilation to the device topology. For each modulus $N$, we focus on detecting a statistically significant excess of probability mass in the theoretically predicted windows around $sL/r$ relative to the uniform baseline. The formulas of the uniform baseline and the empirical probability used for the experiments, rigorously derived in the Appendix A.A, are the following:
$b=\frac{|acceptance\_bin|\times(2\omega_{0}+1)}{L}\text{ and }\hat{p}=\frac{hits}{shots}$
We consider three composite numbers: $N=15$, $N=21$, and $N=35$. In the Appendix subsection A.G, we further discuss the shape of the empirical histograms compared to the ideal noiseless expected distributions.
@@ 5 Limitations and Challenges
Our experiments focus primarily on detecting a statistically significant quantum signal in QPE histograms, rather than fully establishing polylogarithmic scaling of the expected runtime for a fixed acceptance parameter $k$ (see Appendix A.F). For each small modulus $N$, we tune window sizes and shot counts to test whether the mass in acceptance regions exceeds a uniform baseline. This is sufficient to detect coherence in the presence of noise but does not yet constitute a full scalability study.
@@ Acknowledgments
We gratefully acknowledge Evgueni Rousselot for his initial experiments, which were instrumental in cenabling and motivating the investigations further developed in this paper. We also would like to thank Dr. Martin Ekerå for pointing out some theoretical mistakes and taking the time to give his insights.
@@ A.B Per-Run Success Probability
A.C Importance of the Quantum Signal
@@
A.D Certifying Substantial Quantum Signal in Practice
@@
A.E Shor in the Ideal Regime
@@
A.F Shor in the Experimental Regime
@@
A.G Effect of Decoherence on the QPE Output Distribution